OpenStudy (anonymous):

FAN AND MEDAL Suppose that a family has an equally likely chance of having a cat or a dog. If they have two pets, they could have 1 dog and 1 cat, they could have 2 dogs, or they could have 2 cats. What is the theoretical probability that the family has two dogs or two cats? Describe how to use two coins to simulate which two pets the family has.

OpenStudy (welshfella):

yes you can use 2 coins to simulate this the possible results from tossing 2 coins together is HH, HT TH and HH or if you like DD, DC , CD or CC for cats and dogs. and the theoretical probability of each is 1/4 since these 4 are all equally likely . So what do you think the theoretical probability of 2 dogs or 2 cats is? i

OpenStudy (welshfella):

P(DD) = 1/4 , P(CC) = 1/4

OpenStudy (welshfella):

P(CC or DD) = ?

OpenStudy (welshfella):

or you can look at it another way - DD and CC are 2 out of the 4 possible outcomes so ?

OpenStudy (welshfella):

HOLD ON thers a mistake there the order of cat and dog does not matter so there are 3 possibilities DD, DC and CC making the theoretical property of each of these = 1/3 so P(DD or CC) = ?